Session: Feb 2012 (session 1)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Yebes (Pablo de Vicente at Wed Feb 29 12:21:06 2012) Failure: ANT
The observation has failed from 00:02 (scan 0825) to 08:13 (scan 1429) due to an antenna control problem. After 08:13 the observation resumed correctly.

Shanghai (Bo Xia at Thu Mar 01 06:48:44 2012) Success:

Effelsberg (Uwe Bach at Thu Mar 01 09:26:21 2012) Success:

Svetloe (I.Stepanova at Thu Mar 01 11:35:50 2012) Success:
No scans No01039,1040,No1302,1303,1304,1311,1312,1313,1385,1386 - too late on source many times.

Onsala (Michael Lindqvist at Mon Mar 05 10:57:22 2012) Minor failures:
Problem with the IF-level (strange bump in the bandpass), see log for details. Added an aqualizer at around 07:50.

Metsahovi (E. Rastorgueva at Mon Mar 05 11:52:12 2012) Failure: RECO
The experiment used recording at 1Gbps, which our Mark5 defective I/O board fails to support. Recording was not done.

Jodrell_Bank (Alastair Gunn at Mon Mar 19 15:13:46 2012) Success:
No known problems.